The following sample data are from a normal population: 14, 12, 16, 19, 17, 15, 10, 9.

(a)

What is the point estimate of the population mean?

(b)

What is the point estimate of the population standard deviation? (Round your answer to three decimal places.)

(c)

With 95% confidence, what is the margin of error for the estimation of the population mean? (Round your answer to one decimal place.)

(d)

What is the 95% confidence interval for the population mean? (Round your answer to one decimal place.)

to
Answer #1

Solution :

We have,

14, 12, 16, 19, 17, 15, 10, 9.

(a)

. = 112 / 8 = 14

(b)

s = 3.464

(c)

t /2,df = 1.895

Margin of error = E = t/2,df * (s /n)

= 1.895 * (3.464 / 8)

Margin of error = E = 2.3

(d)

The 95% confidence interval estimate of the population mean is,

- E < < + E

14 - 2.3 < < 14 + 2.3

11.7 < < 16.3

(11.7 ,16.3)